### v29

# A short computer program that computes in the limit a non-computable function from $\mathbb{N}$ to $\mathbb{N}$

Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\gamma_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\gamma(n)$. The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation. 2020 Mathematics Subject Classification: 03D20.
1 Predicate $\mathcal{K}$} of the current mathematical knowledge
In sections 14, $\mathcal{K}$} denotes both the predicate satisfied by every currently known theorem and the set of all currently known theorems. The set $\mathcal{K}$} is time-dependent. Observations 15 explain the definition of $\mathcal{K}$}. This definition is intentionally simple, which restricts the set $\mathcal{K}$}. Observations 14 hold at any time.
Observation 1.
Every written down theorem from past or present belongs to $\mathcal{K}$}.
Observation 2.
If $\Delta\in{\mathcal{K}}$, then every particular case of $\Delta$ belongs to $\mathcal{K}$}. For example,
$\{\emptyset\cup\{n\}=\{n\}:~n\in\mathbb{N}\}\subseteq{\mathcal{K}}$
Observation 3.
Since the Replacement Axiom is a scheme of axioms, only finitely many axioms of $ZFC$ belong to $\mathcal{K}$}.
Observation 4.
$\mathcal{K}$} is not deductively closed.
Proof.
There exists $n\in\mathbb{N}$ such that $\neg{\mathcal{K}}(\underbrace{(2\cdot 2=4)\wedge\ldots\wedge(2\cdot 2=4)}_{{\rm conjunction~of~}n+3{\rm~statements}})$. ∎
Observation 5.
Currently,
$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~even}))\wedge$
$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~odd}))$
Hence, $\mathcal{K}$} is not deductively closed now.
2 A program that computes in the limit a function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$
For $n\in\mathbb{N}$, let
$E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$
Theorem 1.
([8, p. 118]). There exists a limit-computable function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$.
We present an alternative proof of Theorem 1. For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$, see [10]. The term "dominated" in the title of [10] means "eventually dominated". Flowchart 1 shows a semi-algorithm which computes $\gamma(n)$ in the limit, see [10].
Flowchart 1
A semi-algorithm which computes $\gamma(n)$ in the limit
Corollary 1.
For every $k\in\mathbb{N}$, the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable.
Proposition 1.
If $k\in\mathbb{N}$, then the statement "the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable" belongs to $\mathcal{K}$}.
Proof.
It follows from Corollary 1 and Observations 1 and 2. ∎
Flowchart 2 shows a simpler semi-algorithm which computes $\gamma(n)$ in the limit.
Flowchart 2
A simpler semi-algorithm which computes $\gamma(n)$ in the limit
Lemma 1.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 2 does not exceed the number printed by Flowchart 1.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 2.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 1 does not exceed the number printed by Flowchart 2.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the system of equations
$\widetilde{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Theorem 2.
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the same number.
Proof.
It follows from Lemmas 1 and 2. ∎
MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if max(op(X[q]))<max(op(X[p])) and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while:
The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$.
For $n\in\mathbb{N}$, $\theta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [10] and Lemma 3 in [9], it follows that the function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $\theta$ in the limit.
3 A program that computes in the limit a function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
The Davis-Putnam-Robinson-Matiyasevich theorem states that every listable set ${\mathcal{M}}\subseteq{\mathbb{N}}^{n}$ $(n\in\mathbb{N}\setminus\{0\})$ has a Diophantine representation, that is
$(a_{1},\ldots,a_{n})\in{\mathcal{M}}\Longleftrightarrow\exists x_{1},\ldots,x_{m}\in\mathbb{N}~~W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ (R)
for some polynomial $W$ with integer coefficients, see [5]. The representation (R) is said to be single-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has at most one solution $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$. The representation (R) is said to be finite-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has only finitely many solutions $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$.
Conjecture 1.
([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
Conjecture 2.
([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a finite-fold Diophantine representation.
Let $\Phi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. For $n\in\mathbb{N}$, let
$g(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Phi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
The function $g:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Phi$ holds. Currently,
$(\neg{\mathcal{K}}(\Phi))\wedge(\neg{\mathcal{K}}(\neg\Phi))\wedge(\neg{\mathcal{K}}(g~is~computable))\wedge(\neg{\mathcal{K}}(g~is~uncomputable))$
Let $\Psi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation. For $n\in\mathbb{N}$, let
$h(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Psi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
The function $h:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Psi$ holds. Currently,
$(\neg{\mathcal{K}}(\Psi))\wedge(\neg{\mathcal{K}}(\neg\Psi))\wedge(\neg{\mathcal{K}}(h~is~computable))\wedge(\neg{\mathcal{K}}(h~is~uncomputable))$
Lemma 3.
The function $g$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The function $h$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation.
Proof.
It follows from Theorem 1. ∎
For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The computability of $\beta$ is unknown.
Theorem 3.
The function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [10]. Flowchart 3 shows a semi-algorithm which computes $\beta(n)$ in the limit, see [10].
Flowchart 3
A semi-algorithm which computes $\beta(n)$ in the limit
Flowchart 4 shows a simpler semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 4
A simpler semi-algorithm which computes $\beta(n)$ in the limit
Lemma 4.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 4 does not exceed the number printed by Flowchart 3.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 5.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 3 does not exceed the number printed by Flowchart 4.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system of equations
$\widehat{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
By this and the inclusion $\widehat{\mathcal{S}}\supseteq{\mathcal{S}}$, $\widehat{\mathcal{S}}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
Theorem 4.
For every $n,m\in\mathbb{N}$, Flowcharts 3 and 4 print the same number.
Proof.
It follows from Lemmas 4 and 5. ∎
The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if q<>p and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while:
Statement 1.
There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
Statement 1 follows constructively from Theorem 3 by taking $f=\beta$ because the following conjunction
$({\neg}{\mathcal{K}}(\beta{\rm~is~computable}))~\wedge~(\neg{\mathcal{K}}(\beta{\rm~is~uncomputable}))$
holds. Statement 1 follows non-constructively from Lemma 3 by taking $f=g$ because the following conjunction
$({\neg}{\mathcal{K}}(g{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(g{\rm~is~uncomputable}))$
holds. ∎
Since the function $\gamma$ in Theorem 1 is not computable, Statement 1 does not follow from Theorem 1.
Statement 2.
In Statement 1, we can require that there exists a computer program which takes as input a non-negative integer $n$ and prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
Proof.
Any computer program that implements the semi-algorithm shown in Flowchart 3 or 4 is correct. ∎
Proposition 2.
Statement 1 strengthens a mathematical theorem. Statement 1 refers to the current mathematical knowledge and may be false in the future. Statement 1 does not express what is currently unproved in mathematics.
Proof.
Statement 1 strengthens Statement 1 without the epistemic condition. The weakened Statement 1 is a theorem which follows from Theorem 1. Statement 1 claims that some mathematically defined function $f:\mathbb{N}\to\mathbb{N}$ satisfies
($f$ is computable in the limit) $\wedge$ (${\neg}{\mathcal{K}}(f$ is computable)) $\wedge$ $({\neg}{\mathcal{K}}(f$ is uncomputable)) $\wedge$
($f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation)
Conjecture 1 disproves Statement 1. ∎
Statement 3.
Statement 1 holds for finite-fold Diophantine representations.
Proof.
It follows from Lemma 3 by taking $f=h$ because the following conjunction
$({\neg}{\mathcal{K}}(h{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(h{\rm~is~uncomputable}))$
holds. ∎
Statement 3 strengthens Statement 1. For Statement 3, there is no known computer program that computes $f$ in the limit.
4 A currently true statement $\bf G$} of the form $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability $\ldots$", where $f$ is conjecturally uncomputable, $\bf G$} remains unproven when $f$ is computable, and $\bf G$} holds forever when $f$ is uncomputable
Let $\bf G$} denote the following statement: $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability such that $f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation and there is a computer program which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
Proposition 3.
The statement $\bf G$} has all properties from the title of this section.
Proposition 4.
The statement $\bf G$} may hold when ${\mathcal{K}}(\beta{\rm~is~computable})$.
Proof.
For $n\in\mathbb{N}$, $\xi(n)$ denotes the smallest $b\in\mathbb{N}\setminus\{0\}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{Q}}^{n+1}$, then $\mathcal{S}$} has a solution which consists of rationals whose numerators and denominators belong to $\{-b,\ldots,b\}$. The function $\xi:\mathbb{N}\to\mathbb{N}$ is computable in the limit. We skip the proof, which is similar to the proof that $\gamma$ is computable in the limit. Currently,
$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))$
At any future time, the conjunction
$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))\wedge{\mathcal{K}}(\beta{\rm~is~computable})$
may hold and implies the statement $\bf G$} with $f=\xi+\beta$. ∎
Summarizing, the predicate $\mathcal{K}$} can strengthen existential mathematical statements when $\mathcal{K}$} is inserted after $\exists$ and refers to a part of the statement.
5 Predicate $\mathcal{K}$} of the written down mathematical knowledge
In this section, $\mathcal{K}$} denotes both the predicate satisfied by every written down theorem and the finite set of all written down theorems. It changes what is taken as known in mathematics.
Proposition 5.
Since $\mathcal{K}$} is finite, there exists $k\in\mathbb{N}$ such that the computability of the function
$\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$
is unknown. For this $k$, Statements 1 and 3 hold when $f(n)=k+\gamma(n)$.
Proposition 5 contradicts Proposition 1 with the right definition of $\mathcal{K}$}.