### v33

# An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$

Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. We prove that no algorithm takes as input a non-negative integer $n$ and decides whether or not $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$$\wedge$ $(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}~(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
2020 Mathematics Subject Classification: 03D20, 11U05.
Key words and phrases: arithmetic of $\mathbb{N}$, computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
1 The Collatz problem leads to a short computer program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
Definition 1.
(cf. [10, pp. 233–235]). A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
By Definition 1, a function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit when there exists an infinite computation which takes as input a non-negative integer $n$ and prints a non-negative integer on each iteration and prints $f(n)$ on each sufficiently high iteration.
It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable, see Theorem 1. Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit. So far, short computer programs can only compute in the limit functions from $\mathbb{N}$ to $\mathbb{N}$ whose computability is proven or unknown.
Lemma 1.
For every $n\in\mathbb{N}$,
$\frac{\textstyle\text{sign}(n-1)\cdot(2n+(1-(-1)^{n})\cdot(5n+2))}{\textstyle 4}=\left\{\begin{array}[]{rl}0{\rm,}&{\rm if~}n=1\\ \frac{\textstyle n}{\textstyle 2}{\rm,}&{\rm if~}n{\rm~is~even}\\ 3n+1{\rm,}&{\rm if~}n{\rm~is~odd~and~}n\neq 1\end{array}\right$.
MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. By Lemma 1, the following program in MuPAD computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$.
input("Input a non-negative integer n",n):
while TRUE do
print(sign(n)):
n:=sign(n-1)*(2*n+(1-(-1)^n)*(5*n+2))/4:
end_while:
The computability of $\gamma$ is unknown, see [1, p. 79]. The Collatz conjecture implies that $\gamma(n)=0$ for every $n\in\mathbb{N}$.
2 A limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
For $n\in\mathbb{N}$, let
$E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$
Theorem 1.
([9, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
We present an alternative proof of Theorem 1. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$, see [12]. The term "dominated" in the title of [12] means "eventually dominated". Flowchart 1 shows a semi-algorithm which computes $f(n)$ in the limit, see [12].
Flowchart 1
A semi-algorithm which computes $f(n)$ in the limit
3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers that concerns the arithmetic of $\mathbb{N}$
Flowchart 2 shows a simpler semi-algorithm which computes $f(n)$ in the limit.
Flowchart 2
A simpler semi-algorithm which computes $f(n)$ in the limit
Lemma 2.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 2 does not exceed the number printed by Flowchart 1.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 3.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 1 does not exceed the number printed by Flowchart 2.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
$\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Theorem 2.
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the same number.
Proof.
It follows from Lemmas 2 and 3. ∎
Definition 2.
An approximation of a tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ is a tuple $(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}$ such that
$(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
Observation 1.
For every $n\in\mathbb{N}$, there exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
${\rm card}({\mathcal{A}}(n))\leqslant 2^{\textstyle{\rm card}(E_{n})}=2^{\textstyle n+1+2\cdot(n+1)^{3}}$
and every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in ${\mathcal{A}}(n)$.
Observation 2.
For every $n\in\mathbb{N}$, $f(n)$ equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
Observation 3.
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\{0,\ldots,m\}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
Theorem 3.
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence: every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,m\}^{n+1}$.
Proof.
Since the function $f$ is not computable, it follows from Observation 2. ∎
4 A short program in MuPAD that computes $f$ in the limit
The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if max(op(X[q]))<max(op(X[p])) and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while:
5 An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $h$ in the limit.
Theorem 4.
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence:
$\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}$
$(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
Proof.
It holds because the function $h$ is not computable. ∎
Corollary 1.
No algorithm takes as input a non-negative integer $n$ and returns the logical value of the following sentence:
$\exists p,q\in\mathbb{N}~((n=2^{p}\cdot 3^{q})\wedge$
$(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
$(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$
6 A limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
The Davis-Putnam-Robinson-Matiyasevich theorem states that every listable set ${\mathcal{M}}\subseteq{\mathbb{N}}^{n}$ $(n\in\mathbb{N}\setminus\{0\})$ has a Diophantine representation, that is
$(a_{1},\ldots,a_{n})\in{\mathcal{M}}\Longleftrightarrow\exists x_{1},\ldots,x_{m}\in\mathbb{N}~~W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ (R)
for some polynomial $W$ with integer coefficients, see [6]. The representation (R) is said to be single-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has at most one solution $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$.
Hypothesis 1.
([2], [3], [4], [5, pp. 341–342], [7, p. 42], [8, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The computability of $\beta$ is unknown.
Theorem 5.
The function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [12]. Flowchart 3 shows a semi-algorithm which computes $\beta(n)$ in the limit, see [12].
Flowchart 3
A semi-algorithm which computes $\beta(n)$ in the limit
7 A short program in MuPAD that computes $\beta$ in the limit
Flowchart 4 shows a simpler semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 4
A simpler semi-algorithm which computes $\beta(n)$ in the limit
Lemma 4.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 4 does not exceed the number printed by Flowchart 3.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 5.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 3 does not exceed the number printed by Flowchart 4.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{\mathcal{S}}$, where
$\widehat{\mathcal{S}}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
By this and the inclusion $\widehat{\mathcal{S}}\supseteq{\mathcal{S}}$, $\widehat{\mathcal{S}}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
Theorem 6.
For every $n,m\in\mathbb{N}$, Flowcharts 3 and 4 print the same number.
Proof.
It follows from Lemmas 4 and 5. ∎
The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if q<>p and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while: